### v1

# references.bib

Abstract
This paper studies the linear system identification problem in the general case where the disturbance is sub-Gaussian, correlated, and possibly adversarial. First, we consider the case with noncentral (nonzero-mean) disturbances for which the ordinary least-squares (OLS) method fails to correctly identify the system. We prove that the $l_{1}$-norm estimator accurately identifies the system under the condition that each disturbance has equal probabilities of being positive or negative. This condition restricts the sign of each disturbance but allows its magnitude to be arbitrary. Second, we consider the case where each disturbance is adversarial with the model that the attack times happen occasionally but the distributions of the attack values are completely arbitrary. We show that when the probability of having an attack at a given time is less than 0.5, the $l_{1}$-norm estimator prevails against any adversarial noncentral disturbances and the exact recovery is achieved within a finite time. These results pave the way to effectively defend against arbitrarily large noncentral attacks in safety-critical systems.
I Introduction
The system identification aims to identify the unknown parameters that govern an underlying dynamical system, given the history of states determined by the true parameters and the disturbances. In real-world applications, one may not be able to model the system accurately, and therefore it is vital to use the collected online data to identify the system. The identification is often followed by the control of the system, and a small identification error improves the quality of the control and affects the stability and optimality of adaptive control strategies [kumar1986stochastic, ioannou2012robust]. In this paper, we focus on the linear system identification, where the states are determined by $x_{t+1}=A^{*}x_{t}+w_{t},~{}t=0,\dots,T-1$, and the goal is to find the true unknown matrix $A^{*}$ using the state trajectory, despite the presence of disturbances $w_{t}$.
The challenge of this problem is that the states are correlated with each other, even when the disturbances are assumed to be independent and identically distributed (i.i.d.). Early papers have shown that the ordinary least-squares (OLS) estimator $\operatorname*{arg\,min}_{A\in\mathbb{R}^{d\times d}}\sum_{t=0}^{T-1}\|x_{t+1}-Ax_{t}\|_{2}^{2}$ overcomes this correlation and achieve the consistency and the convergence towards the true matrix [ljung1976ols, moore1978strong, lai1982least]. Recently, instead of the asymptotic properties, several studies have focused on non-asymptotic analyses and provided an error rate of $O\bigr{(}\frac{1}{\sqrt{T}}\bigr{)}$ between the true matrix and OLS estimates under i.i.d. zero-mean Gaussian and sub-Gaussian disturbances after finite samples are collected [simchowitz2018learningwithout, jedra2020finite, sarkar2021finite]. These results were extended to use the OLS estimates for the linear quadratic control with unknown matrices, in a way that once a small enough estimation error is attained, the estimates are treated as the true matrices, achieving $O(\sqrt{T})$ regret [mania2019certainty, simchowitz2020naive]. While OLS is indeed a statistically optimal estimator in such i.i.d. cases, little is known about the system identification when the disturbances are correlated or possibly selected adversarially based on the past information.
Recently, in the machine learning literature, online control problem with any bounded adversarial disturbance, called online nonstochastic control, has gained considerable attention, among which [hazan2020nonstochastic, simchowitz2020improper, chen2021blackbox] leveraged OLS method to recover the matrix and attained $O(T^{2/3})$ regret. However, this polynomial regret implies a substantial cumulative loss, leading to an impractical algorithm for real-world applications. To address more applicable situations, [feng2021adv, yalcin2023exact, zhang2024exact] assumed that adversarial attacks occur with probability less than 1, and exactly recover the true matrix with a non-smooth $l_{2}$-norm estimator defined by $\operatorname*{arg\,min}_{A\in\mathbb{R}^{d\times d}}\sum_{t=0}^{T-1}\|x_{t+1}-Ax_{t}\|_{2}$. In this case, practicality is achieved since the regret does not grow after the exact recovery. However, they assumed that each nonzero adversarial attack follows a zero-mean sub-Gaussian distribution, which restricts the range of possible attacks.
In this paper, we allow noncentral (nonzero-mean) sub-Gaussian adversarial disturbances and use the $l_{1}$-norm estimator $\operatorname*{arg\,min}_{A\in\mathbb{R}^{d\times d}}\sum_{t=0}^{T-1}\|x_{t+1}-Ax_{t}\|_{1}$ to identify the unknown matrix within a finite time. Our work presents the first result in the literature for the exact retrieval of the true matrix even in the presence of adversarial noncentral disturbances. Under the assumption that the probability of having an attack at a given time is $p<1$, we construct two scenarios of the adversarial disturbance structure:
  1. 1.
    symmetric disturbance around zero scaled by an amplifying random vector (see Assumption 5) with $p<1$,
  2. 2.
    any adversarial noncentral disturbance with $p<0.5$.
The first scenario is applicable to the domain where the positive and negative disturbances occur with equal chance, each of which can be scaled arbitrarily. For example, beliefs based on behavioral bias may cause fluctuations in the stock market [lo2004adaptive] or energy demand [werth2021energy], which can be explained by disturbances of arbitrary magnitude. A more practical situation arises in the second scenario: it applies when an extremely large attack with a nonzero mean occasionally affects the system, such as natural disaster on power systems [buldyrev2010cat, wang2016power, yan2020power, trakas2022st], unanticipated malicious cyber attacks [pasqualetti2013cyber, duo2022cyber], and others. In this paper, we will show that the $l_{1}$-norm estimator is indeed a robust estimator capable of successfully prevailing over such disturbances. Note that while the paper focuses on autonomous linear systems, it can be readily generalized to linear systems with inputs as well as nonlinear systems linearly parameterized by basis functions. This generalization can be achieved by adopting the formulation and technique studied in [zhang2024exact], but this will be left as future work due to space restrictions.
The paper is organized as follows. In Sections II and III, we introduce the preliminaries and formulate the problem, respectively. In Section IV, we study the two scenarios of adversarial noncentral disturbances and show that the exact recovery is achieved within finite time. In Section V, we present numerical experiments that support our main results. Finally, concluding remarks are provided in Section VI.
Notation. Let $\mathbb{R}^{d}$ denote the set of $n$-dimensional vector and $\mathbb{R}^{d\times d}$ denote the set of $d\times d$ matrices. For a matrix $A$, $\|A\|_{2}$ denotes the operator norm of the matrix. For a vector $x$, $\|x\|_{1}$ denotes the $l_{1}$-norm of the vector and $\|x\|_{1}$ denotes the $l_{2}$-norm of the vector. $x^{T}$ denotes the tranpose of the vector $x$. $x>0$ means that all entries of $x$ are positive. For two vectors $x$ and $y$ of the same dimensions, let $x\circ y$ denote entry-wise product of the same dimensions as $x$ and $y$. Let $\mathbb{E}$ denote the expectation operator. For the event $\mathcal{E}$, let $\mathbb{P}(\mathcal{E})$ denote the probability of the event. Also, let $\mathbb{I}(\mathcal{E})$ denote the indicator of the event, meaning that $\mathbb{I}(\mathcal{E})=1$ if the event occurs and $\mathbb{I}(\mathcal{E})=0$ otherwise. We use $O(\cdot)$ for the big-O notation.
II Preliminaries
In this work, we consider each disturbance vector to follow sub-Gaussian distribution. We use the definition of sub-Gaussian given in [vershynin2018high].
Definition 1 (sub-Gaussian scalar variables)
A random variable $w\in\mathbb{R}$ is called sub-Gaussian if there exists $c>0$ such that
$\mathbb{E}\Bigr{[}\exp\Bigr{(}\frac{w^{2}}{c^{2}}\Bigr{)}\Bigr{]}\leq 2.$ (1)
Its sub-Gaussian norm is denoted by $\|w\|_{\psi_{2}}$ and defined as
$\|w\|_{\psi_{2}}=\inf\left\{c>0:\mathbb{E}\Bigr{[}\exp\Bigr{(}\frac{w^{2}}{c^{2}}\Bigr{)}\Bigr{]}\leq 2\right\}.$ (2)
Note that the $\psi_{2}$-norm satisfies all properties of norm, including positive definiteness, homogeneity, and triangle inequality. We introduce the equivalent definition of sub-Gaussian variables in the case where $\mathbb{E}[w]=0$.
Definition 2 (sub-Gaussian scalar variables (MGF))
A random variable $w\in\mathbb{R}$ with $\mathbb{E}[w]=0$ is called sub-Gaussian if there exists $\eta_{w}>0$ such that the moment generating function (MGF) of $w$ satisfies
$\mathbb{E}[\exp(\lambda w)]\leq\exp(\lambda^{2}\eta_{w}^{2})$ (3)
for all $\lambda\in\mathbb{R}$. For any sub-Gaussian variable that satisfies (1) and $\mathbb{E}[w]=0$, (3) is satisfied with $\eta_{w}\leq 6\sqrt{e}\|w\|_{\psi_{2}}$ [vershynin2018high].
To leverage the equivalent definition (3) only holding for zero-mean random variables, the centering lemma in [vershynin2018high] is useful. We specify the previously omitted absolute constant in the lemma below.
Lemma 1 (Centering lemma, [vershynin2018high])
If $w$ is a sub-Gaussian variable satisfying (1), then $w-\mathbb{E}[w]$ is also sub-Gaussian and it holds that
$\|w-\mathbb{E}[w]\|_{\psi_{2}}\leq\left(1+\sqrt{\frac{\pi}{\log 2}}\right)\|w\|_{\psi_{2}}.$ (4)
Real-world systems have a limit on their actuators so they cannot accept arbitrarily large inputs. This enables us to assume that each disturbance follows sub-Gaussian variables of which the tail event rarely occurs. Hoeffding’s inequality provides sharp bounds on the tail event.
Lemma 2 (Hoeffding’s inequality, [vershynin2018high])
Suppose that a random variable $w$ satisfies (3) with the parameter $\eta_{w}$. Then, for all $s>0$, we have
$\mathbb{P}(w\geq s)\leq\exp\Bigr{(}-\frac{s^{2}}{(2\eta_{w})^{2}}\Bigr{)},$ (5)
$\mathbb{P}(w\leq-s)\leq\exp\Bigr{(}-\frac{s^{2}}{(2\eta_{w})^{2}}\Bigr{)}.$ (6)
To provide the analysis of high-dimensional systems, we provide the definition of multi-dimensional sub-Gaussian variables given in [vershynin2018high].
Definition 3 (sub-Gaussian vector variables)
A random vector $w\in\mathbb{R}^{d}$ is called sub-Gaussian if for every $x\in\mathbb{R}^{d}$, $w^{T}x$ is a sub-Gaussian scalar variable defined in Definition 1. Its norm is defined as
$\|w\|_{\psi_{2}}=\sup_{\|x\|_{2}\leq 1,x\in\mathbb{R}^{d}}\|w^{T}x\|_{\psi_{2}}.$ (7)
Now, we provide the subdifferential of the $l_{1}$-norm used for the analysis of the $l_{1}$-norm estimator.
Definition 4 (Subdifferential of the $l_{1}$-norm)
Given a vector $x\in\mathbb{R}^{d}$ and its entries $x^{1},\dots,x^{d}$, the subdifferential of $\|x\|_{1}$ is denoted by $\partial\|x\|_{1}$ and defined as $[\text{sgn}(x^{1}),\dots,\text{sgn}(x^{d})]^{T}$ where
$\text{sgn}(z)=\begin{dcases*}1,&if $z>0$,\\ [-1,1],&if $z=0$,\\ -1,&if $z<0$.\end{dcases*}$ (8)
The decomposability of the subdifferential along the coordinates simplifies the analysis.
III Problem Formulation
Consider a general linear time-invariant dynamical system with additive disturbance
$x_{t+1}=A^{*}x_{t}+w_{t},\quad t=0,1,\dots,T-1,$ (9)
where $A^{*}\in\mathbb{R}^{d\times d}$ is the unknown true matrix, $x_{t}\in\mathbb{R}^{d}$ is the state, and $w_{t}\in\mathbb{R}^{d}$ is a disturbance injected into the system at time $t$. Our goal is to identify the true matrix $A^{*}$, given the state trajectory $x_{0},\dots,x_{T}$. We assume that $A^{*}$ has bounded operator norm and each $x_{0},w_{0},\dots,w_{T-1}$ is sub-Gaussian to prevent an unbounded growth of the system. However, we allow adversarial noncentral disturbances, meaning that an adversary can design the disturbance $w_{t}$ based on the previous information $\mathcal{F}_{t}=\{x_{0},w_{0},\dots,w_{t-1}\}$, and $\mathbb{E}[w_{t}~{}|~{}\mathcal{F}_{t}]$ is not necessarily zero. We formally present the relevant assumptions below.
Assumption 1 (Operator norm)
It holds that $\|A^{*}\|_{2}<1$ (this condition is somewhat stronger than stability).
Assumption 2 (Maximum sub-Gaussian norm)
There exists $\sigma_{w}>0$ such that $\|x_{0}\|_{\psi_{2}}\leq\sigma_{w}$ and $\|w_{t}\|_{\psi_{2}}\leq\sigma_{w}$ conditioned on $\mathcal{F}_{t}$ for all $t\geq 0$ and for all $\mathcal{F}_{t}$.
Our goal is to achieve a finite-time exact recovery, meaning that the error between the true matrix and the estimate should be exactly zero after a finite time. However, it is well-established that the lower bound of the error is $O\bigr{(}\frac{1}{\sqrt{T}}\bigr{)}$ if disturbances follow i.i.d. zero-mean Gaussian distribution [simchowitz2018learningwithout]. Our disturbance structure includes this special distribution, and therefore the lower bound $O\bigr{(}\frac{1}{\sqrt{T}}\bigr{)}$ does not allow an exact recovery without a further assumption also adopted by [feng2021adv, yalcin2023exact, zhang2024exact].
Assumption 3 (Probabilistic Attack)
Let $w_{t}^{i}$ denote the $i^{\text{th}}$ entry of the disturbance $w_{t}$. Each coordinate of the disturbance has a positive probability of being zero conditioned on $\mathcal{F}_{t}$, meaning that there exists $p<1$ such that
$\mathbb{P}(w_{t}^{i}=0~{}|~{}\mathcal{F}_{t})\geq 1-p$ (10)
holds for all $i=1,\dots,d$, as well as $t$ and $\mathcal{F}_{t}$.
Note that Assumption 3 implies that the system is not under attack at time $t$ with probability at least $1-p$. Now, we consider the following $l_{1}$-norm estimator at time $T$:
$\min_{A\in\mathbb{R}^{d\times d}}\sum_{t=0}^{T-1}\|x_{t+1}-Ax_{t}\|_{1},$ (11)
given a state trajectory $x_{0},\dots,x_{T}$. We will show that $l_{1}$-estimator successfully overcomes noncentral distributions and achieves a finite-time exact recovery.
IV Main Results
In this section, under the proper assumptions, we will show that $A^{*}$ is a solution to the $l_{1}$-norm estimator after a finite time. Let $\hat{A}_{T}$ denote any estimate obtained from (11), which can be equivalently written as
$\hat{A}_{T}\in\operatorname*{arg\,min}_{A\in\mathbb{R}^{d\times d}}\sum_{t=0}^{T-1}\|(A^{*}-A)x_{t}+w_{t}\|_{1}$ (12)
due to the system dynamics (9). In the next subsection, we first elucidate the exact recovery conditions of the $l_{1}$-norm estimator.
IV-A Conditions for the Exact Recovery
In this subsection, we provide necessary and sufficient conditions for the exact recovery. For the following theorem, let $x_{t}^{i}$ and $w_{t}^{i}$ denote the $i^{\text{th}}$ entries of $x_{t}$ and $w_{t}$, respectively. Define the random variables $z_{t}^{ij+}$ and $z_{t}^{ij-}$ as follows:
[…]
Theorem 1
$A^{*}$ is a solution to the $l_{1}$-norm estimator (11) at time $T$ if and only if
$\sum_{t=0}^{T-1}z_{t}^{ij+}\geq 0,$ (14a)
$\sum_{t=0}^{T-1}z_{t}^{ij-}\geq 0$ (14b)
for all $i,j\in\{1,\dots,d\}$.
Proof:
Since the non-differential minimization problem (12) is convex and unconstrained, the Karush-Kuhn-Tucker (KKT) conditions provide the necessary and sufficient conditions for optimality:
$0\in\sum_{t=0}^{T-1}\partial\|(A^{*}-\hat{A}_{T})x_{t}+w_{t}\|_{1}\cdot x_{t}^{T}.$ (15)
Thus, $A^{*}$ is a solution to (12) (or equivalently (11)) if and only if
$0\in\sum_{t=0}^{T-1}\partial\|w_{t}\|_{1}\cdot x_{t}^{T}.$ (16)
Note that both sides of (16) are $d\times d$ matrices. Using Definition 4, the equivalent form of (16) is
$-\sum_{t:w_{t}^{i}=0}|x_{t}^{j}|\leq-\sum_{t:w_{t}^{i}<0}x_{t}^{j}+\sum_{t:w_{t}^{i}>0}x_{t}^{j}\leq\sum_{t:w_{t}^{i}=0}|x_{t}^{j}|$ (17)
for all $i,j\in\{1,\dots,d\}$. One can separate (17) into two inequalities and attain $2d^{2}$ number of inequalities, which are again rearranged to (14a) and (14b). ∎
To attain the exact recovery, the inequalities in (14) must hold. Since the left-hand sides of (14a) and (14b) are random variables, it is necessary to study their expected values and ensure that they are sufficiently positive with a high probability. We require a mild assumption stating that no adversarial disturbance can deceive the system to the extent that each coordinate of the next state collapses to absolute zero, which ensures sufficient exploration of the state trajectory.
Assumption 4 (Non-degeneracy)
There exists $\lambda>0$ such that
$\mathbb{E}[|x_{0}^{i}|]\geq\lambda,\quad$ $i=1,\dots,d,$ (18a)
$\mathbb{E}[|x_{t+1}^{i}|~{}|~{}\mathcal{F}_{t}]\geq\lambda,\quad$ $i=1,\dots,d$ (18b)
for all $t$ and $\mathcal{F}_{t}$.
In the following two subsections, each will present a distinct scenario in which the $l_{1}$-norm estimator achieves the exact recovery despite the presence of adversarial noncentral disturbances.
IV-B Exact Recovery under Sign-restricted Disturbances
In this subsection, we will demonstrate that the exact recovery is achieved under the condition that each disturbance has equal chance of being positive or negative. The structure of the disturbance $w_{t}$ consists of $\alpha_{t}$ and $\beta_{t}$, where $\alpha_{t}$ is a sub-Gaussian vector symmetric around zero and $\beta_{t}>0$ is an amplifying factor. The formal assumption is given below.
Assumption 5 (Sign-restricted disturbances)
Given the filtration $\mathcal{F}_{t}=\{x_{0},w_{0},\dots,w_{t-1}\}$, a disturbance $w_{t}\in\mathbb{R}^{d}$ is formulated as
$w_{t}=\alpha_{t}\circ\beta_{t},$ (19)
where $\alpha_{t}=[\alpha_{t}^{1},\dots,\alpha_{t}^{d}]^{T}$ and $\beta_{t}=[\beta_{t}^{1},\dots,\beta_{t}^{d}]^{T}$ satisfy the following conditions:
  • $\alpha_{t}$ is a sub-Gaussian vector variable (see Definition 3) symmetric around zero conditioned on $\mathcal{F}_{t}$, meaning that for all $i=1,\dots,d$, we have
    $\mathbb{P}(\alpha_{t}^{i}>0~{}|~{}\mathcal{F}_{t})=\mathbb{P}(\alpha_{t}^{i}<0~{}|~{}\mathcal{F}_{t}).$ (20)
  • $\beta_{t}$ is an amplifying random vector conditioned on $\mathcal{F}_{t}$, where for all $i=1,\dots,d$, there exists $B>0$ such that $0<\beta_{t}^{i}\leq B$.
Note that the resulting $w_{t}$ is indeed sub-Gaussian conditioned on $\mathcal{F}_{t}$ since $\alpha_{t}$ is sub-Gaussian and $\beta_{t}$ consists of finite entries.
Remark 1 (Necessity of symmetry assumption)
To formulate adversarial noncentral disturbances, we have introduced $\alpha_{t}$ that is symmetric around zero. We adopted this assumption since the system cannot defend the attack if the adversary has a full information set $\mathcal{F}_{t}$ and the freedom of choosing any nonzero-mean attack. For example, consider a scalar system $x_{t+1}=a^{*}x_{t}+w_{t}$, where the true system satisfies $10<a^{*}<$ 1 and $10<x_{0}<$ 1. Suppose that $w_{t}$ conditioned on $\mathcal{F}_{t}$ is $-\text{sgn}(x_{t})$ for all $t=0,1,\dots,$ where $\text{sgn}(\cdot)$ is defined in (8). In such a case, one always attains
$0<x_{t}<1~{}$ $\text{for}~{}t=0,2,4,\dots,$
$-1<x_{t}<0~{}$ $\text{for}~{}t=1,3,5,\dots.$
However, any valid estimation method tries to minimize the distance between $x_{t+1}$ and $\hat{a}x_{t}$ to find an estimate $\hat{a}$. One can thus only arrive at a negative estimate $\hat{a}$ despite the true system $a^{*}>0$, since we have $\text{sgn}(x_{t})\neq\text{sgn}(x_{t+1})$ for all $t=0,1,\dots$. This occurs since the attacker can always deceive the system if the attack can be universally positive or negative depending on the filtration. The symmetry assumption (20) on $\alpha_{t}$ prevents such a phenomenon and enables system identification.
The disturbance structure (19) includes zero-mean Gaussian random variables since $\alpha_{t}$ can be zero-mean Gaussian distribution and $\beta_{t}$ can be a constant random vector. Our formulation generalizes this by converting a symmetric distribution into a noncentral distribution with a random vector $0<\beta_{t}<\infty$.
Now, we analyze the conditions for the exact recovery shown in Theorem 1. The following lemma provides the expectation information of (14).
Lemma 3
Under Assumptions 3, 4, and 5, we have
$\mathbb{E}\biggr{[}\sum_{t=0}^{T-1}z_{t}^{ij+}\biggr{]}\geq(1-p)\lambda T,$ (21a)
$\mathbb{E}\biggr{[}\sum_{t=0}^{T-1}z_{t}^{ij-}\biggr{]}\geq(1-p)\lambda T.$ (21b)
Proof:
Note that the symmetry assumption (20) implies that the adversarial disturbance $w_{t}$ also satisfies
$\mathbb{P}(w_{t}^{i}>0~{}|~{}\mathcal{F}_{t})=\mathbb{P}(w_{t}^{i}<0~{}|~{}\mathcal{F}_{t})$
for all $\mathcal{F}_{t}$ due to $\beta_{t}>0$. By the definition of $z_{t}^{ij+}$ given in (13), we have
$\mathbb{E}$ $\Bigr{[}\sum_{t=0}^{T-1}z_{t}^{ij+}\Bigr{]}=\sum_{t=0}^{T-1}\mathbb{E}[z_{t}^{ij+}]$
$=\sum_{t=0}^{T-1}\mathbb{E}\bigr{[}|x_{t}^{j}|\cdot\mathbb{I}(w_{t}^{i}=0)+x_{t}^{j}\cdot\mathbb{I}(w_{t}^{i}<0)-x_{t}^{j}\cdot\mathbb{I}(w_{t}^{i}>0)\bigr{]}$
$=\sum_{t=0}^{T-1}\mathbb{E}\Bigr{[}\mathbb{E}\bigr{[}|x_{t}^{j}|\cdot\mathbb{I}(w_{t}^{i}=0)$
$\hskip 56.9055pt+x_{t}^{j}\cdot\mathbb{I}(w_{t}^{i}<0)-x_{t}^{j}\cdot\mathbb{I}(w_{t}^{i}>0)~{}\bigr{|}~{}\mathcal{F}_{t}]\Bigr{]}$
$=\sum_{t=0}^{T-1}\mathbb{E}\Bigr{[}|x_{t}^{j}|\cdot\mathbb{P}(w_{t}^{i}=0~{}|~{}\mathcal{F}_{t})$
$\hskip 33.57425pt+x_{t}^{j}\cdot[\mathbb{P}(w_{t}^{i}<0~{}|~{}\mathcal{F}_{t})-\mathbb{P}(w_{t}^{i}>0~{}|~{}\mathcal{F}_{t})]\Bigr{]}$ (22)
$\geq\sum_{t=0}^{T-1}(1-p)\cdot\mathbb{E}\bigr{[}|x_{t}^{j}|\bigr{]}+0\geq(1-p)\lambda T,$
due to the tower rule and the fact that the filtration $\mathcal{F}_{t}=\{x_{0},w_{0},\dots,w_{t-1}\}$ contains the information of $x_{t}$. The last inequality is due to Assumptions 3 and 4. Similarly, $\mathbb{E}\bigr{[}\sum_{t=0}^{T-1}z_{t}^{ij-}\bigr{]}$ has the same lower bound. ∎
Now, it remains to bound the tail probability of the left-hand sides of (14a) and (14b), for which we will study the sub-Gaussian norms.
Lemma 4
Under Assumptions 1 and 2, we have
$\mathbb{E}\biggr{[}\exp\biggr{(}\lambda\Bigr{[}\sum_{t=0}^{T-1}z_{t}^{ij+}-\mathbb{E}\bigr{[}\sum_{t=0}^{T-1}z_{t}^{ij+}\bigr{]}\Bigr{]}\biggr{)}\biggr{]}$
$\hskip 85.35826pt\leq\exp\biggr{(}\lambda^{2}T\Bigr{(}\frac{\bar{C}\sigma_{w}}{1-\|A^{*}\|_{2}}\Bigr{)}^{2}\biggr{)},$ (23a)
$\mathbb{E}\biggr{[}\exp\biggr{(}\lambda\Bigr{[}\sum_{t=0}^{T-1}z_{t}^{ij-}-\mathbb{E}\bigr{[}\sum_{t=0}^{T-1}z_{t}^{ij-}\bigr{]}\Bigr{]}\biggr{)}\biggr{]}$
$\hskip 85.35826pt\leq\exp\biggr{(}\lambda^{2}T\Bigr{(}\frac{\bar{C}\sigma_{w}}{1-\|A^{*}\|_{2}}\Bigr{)}^{2}\biggr{)}$ (23b)
for all $\lambda\in\mathbb{R}$, where $\bar{C}=6\sqrt{e}\Bigr{(}1+\sqrt{\frac{\pi}{\log 2}}\Bigr{)}$.
Proof:
Recall the definition (7) of the $\psi_{2}$-norm of sub-Gaussian vectors. For a matrix $A\in\mathbb{R}^{d\times d}$ and a vector $x\in\mathbb{R}^{d}$, we have
$\|Ax\|_{\psi_{2}}$ $=\sup_{\|y\|_{2}\leq 1,y\in\mathbb{R}^{d}}\|x^{T}A^{T}y\|_{\psi_{2}}$
$\leq\sup_{\|A^{T}y\|_{2}\leq\|A\|_{2},y\in\mathbb{R}^{d}}\|x^{T}A^{T}y\|_{\psi_{2}}$
$\leq\sup_{\|z\|_{2}\leq\|A\|_{2},z\in\mathbb{R}^{d}}\|x^{T}z\|_{\psi_{2}}$
$=\sup_{\|\tilde{z}\|_{2}\leq 1,\tilde{z}\in\mathbb{R}^{d}}\Bigr{\|}\|A\|_{2}\cdot x^{T}\tilde{z}\Bigr{\|}_{\psi_{2}}=\|A\|_{2}\cdot\|x\|_{\psi_{2}},$ (24)
where the first inequality comes from the relationship
$\|y\|_{2}\leq 1~{}\Rightarrow~{}\|A^{T}y\|_{2}\leq\|A\|_{2}\|y\|_{2}\leq\|A\|_{2}$
and the last equality is by the homogeneity of the $\psi_{2}$-norm.
Now, due to the system dynamics (9), we have
$x_{t}=(A^{*})^{t}x_{0}+\sum_{i=0}^{t-1}(A^{*})^{t-i-1}w_{i},$ (25)
for $t=0,1,\dots,T$. Then, we obtain
$\|x_{t}\|_{\psi_{2}}$ $=\Bigr{\|}(A^{*})^{t}x_{0}+\sum_{i=0}^{t-1}(A^{*})^{t-i-1}w_{i}\Bigr{\|}_{\psi_{2}}$
$\leq\|(A^{*})^{t}x_{0}\|_{\psi_{2}}+\sum_{i=0}^{t-1}\|(A^{*})^{t-i-1}w_{i}\|_{\psi_{2}}$
$\leq\|(A^{*})^{t}\|_{2}\cdot\|x_{0}\|_{\psi_{2}}+\sum_{i=0}^{t-1}\|(A^{*})^{t-i-1}\|_{2}\cdot\|w_{i}\|_{\psi_{2}}$
$\leq\|A^{*}\|_{2}^{t}\cdot\|x_{0}\|_{\psi_{2}}+\sum_{i=0}^{t-1}\|A^{*}\|_{2}^{t-i-1}\cdot\|w_{i}\|_{\psi_{2}}$
$\leq\sigma_{w}\sum_{i=0}^{t}\|A^{*}\|_{2}^{i}\leq\frac{\sigma_{w}}{1-\|A^{*}\|_{2}},$
where the first two inequalities are due to the triangle inequality and (24), respectively, and the last two inequalities come from Assumptions 2 and 1, respectively. Considering standard basis vectors in $\mathbb{R}^{d}$, the definition (7) states that the $\psi_{2}$-norms of one-dimensional projections satisfy $\|x_{t}^{j}\|_{\psi_{2}}\leq\|x_{t}\|_{\psi_{2}}\leq\frac{\sigma_{w}}{1-\|A^{*}\|_{2}}$ for $j=1,\dots,d.$ Since the definition (2) involves the squared variable, alternating the sign or taking the absolute value does not affect the norm of sub-Gaussian variables. Thus, given the filtration $\mathcal{F}_{t}$, we have
$\|z_{t}^{ij+}\|_{\psi_{2}}\leq\frac{\sigma_{w}}{1-\|A^{*}\|_{2}},\quad\|z_{t}^{ij-}\|_{\psi_{2}}\leq\frac{\sigma_{w}}{1-\|A^{*}\|_{2}}$ (26)
conditioned on $\mathcal{F}_{t}$ for all $i,j\in\{1,\dots,d\}$ and all $t$. Using Lemma 1 in turn, one arrives at
$\|z_{t}^{ij+}-\mathbb{E}[z_{t}^{ij+}]\|_{\psi_{2}}\leq\Bigr{(}1+\sqrt{\frac{\pi}{\log 2}}\Bigr{)}\frac{\sigma_{w}}{1-\|A^{*}\|_{2}}$ (27)
conditioned on $\mathcal{F}_{t}$ for all $i,j\in\{1,\dots,d\}$ and all $t$. The same inequality also holds for $z_{t}^{ij-}$. Notice that the expectation of $z_{t}^{ij+}-\mathbb{E}[z_{t}^{ij+}]$ is zero due to the tower rule and we can use the equivalent Definition 2 to obtain
[…]
where $\bar{C}=6\sqrt{e}\Bigr{(}1+\sqrt{\frac{\pi}{\log 2}}\Bigr{)}$. Then, we have
$\mathbb{E}\biggr{[}\exp\Bigr{(}\lambda\Bigr{[}\sum_{t=0}^{T-1}z_{t}^{ij+}-\mathbb{E}\bigr{[}\sum_{t=0}^{T-1}z_{t}^{ij+}\bigr{]}\Bigr{]}\Bigr{)}\biggr{]}$
$=\mathbb{E}\biggr{[}\mathbb{E}\Bigr{[}\exp\Bigr{(}\lambda\Bigr{[}\sum_{t=0}^{T-1}z_{t}^{ij+}-\mathbb{E}\bigr{[}\sum_{t=0}^{T-1}z_{t}^{ij+}\bigr{]}\Bigr{]}\Bigr{)}~{}\Bigr{|}~{}\mathcal{F}_{T-1}\Bigr{]}\biggr{]}$
$=\mathbb{E}\biggr{[}\Bigr{(}\sum_{t=0}^{T-2}z_{t}^{ij+}-\mathbb{E}\bigr{[}\sum_{t=0}^{T-2}z_{t}^{ij+}\bigr{]}\Bigr{)}$
$\hskip 56.9055pt\cdot\mathbb{E}\Bigr{[}\exp\Bigr{(}\lambda\Bigr{[}z_{T-1}^{ij+}-\mathbb{E}\bigr{[}z_{T-1}^{ij+}\bigr{]}\Bigr{]}\Bigr{)}~{}\Bigr{|}~{}\mathcal{F}_{T-1}\Bigr{]}\biggr{]}$
$\leq\mathbb{E}\biggr{[}\Bigr{(}\sum_{t=0}^{T-2}z_{t}^{ij+}-\mathbb{E}\bigr{[}\sum_{t=0}^{T-2}z_{t}^{ij+}\bigr{]}\Bigr{)}\biggr{]}\exp\biggr{(}\lambda^{2}\Bigr{(}\frac{\bar{C}\sigma_{w}}{1-\|A^{*}\|_{2}}\Bigr{)}^{2}\biggr{)}$
$\leq\dots\leq\exp\biggr{(}\lambda^{2}T\Bigr{(}\frac{\bar{C}\sigma_{w}}{1-\|A^{*}\|_{2}}\Bigr{)}^{2}\biggr{)}$
for all $i,j\in\{1,\dots,d\}$, by repetitively applying the tower rule ($z_{t-1}^{ij+}$ is a function of $\mathcal{F}_{t}$) and (28). Due to (26), the same inequality also holds even when $z_{t}^{ij+}$ is replaced by $z_{t}^{ij-}$. ∎
Now, we present one of our main theorems built on Lemmas 3 and 4.
Theorem 2
Suppose that Assumptions 1, 2, 3, 4, and 5 hold. Given $\delta\in(0,1],$ the $l_{1}$-norm estimator defined in (11) achieves the exact recovery after
$T\geq O\biggr{(}\biggr{(}\frac{\sigma_{w}}{(1-p)\lambda(1-\|A^{*}\|_{2})}\biggr{)}^{2}\biggr{)}\cdot\log\Bigr{(}\frac{2d^{2}}{\delta}\Bigr{)}$ (29)
with probability at least $1-\delta$.
Proof:
By Lemma 4, $\sum_{t=0}^{T-1}z_{t}^{ij+}-\mathbb{E}\bigr{[}\sum_{t=0}^{T-1}z_{t}^{ij+}\bigr{]}$ satisfies (3) with the parameter $\sqrt{T}\bigr{(}\frac{\bar{C}\sigma_{w}}{1-\|A^{*}\|_{2}}\bigr{)}$ for all $i,j\in\{1,\dots,d\}$. Then, (6) in Lemma 2 implies that for all $s>0$, we have
$\mathbb{P}\Bigr{(}\sum_{t=0}^{T-1}z_{t}^{ij+}-\mathbb{E}\bigr{[}\sum_{t=0}^{T-1}z_{t}^{ij+}\bigr{]}>-s\Bigr{)}$
$\hskip 85.35826pt\geq 1-\exp\biggr{(}-\frac{s^{2}}{4T\bigr{(}\frac{\bar{C}\sigma_{w}}{1-\|A^{*}\|_{2}}\bigr{)}^{2}}\biggr{)}.$
Since the positive expectation is guaranteed by Lemma 3, one can substitute $\frac{\mathbb{E}[\sum_{t=0}^{T-1}z_{t}^{ij+}]}{2}>0$ into $s$ to obtain
$\mathbb{P}\Bigr{(}\sum_{t=0}^{T-1}z_{t}^{ij+}\geq 0\Bigr{)}$ $\geq\mathbb{P}\Bigr{(}\sum_{t=0}^{T-1}z_{t}^{ij+}>\frac{\mathbb{E}\bigr{[}\sum_{t=0}^{T-1}z_{t}^{ij+}\bigr{]}}{2}\Bigr{)}$
$\geq 1-\exp\biggr{(}-\frac{\mathbb{E}\bigr{[}\sum_{t=0}^{T-1}z_{t}^{ij+}\bigr{]}^{2}}{16T\bigr{(}\frac{\bar{C}\sigma_{w}}{1-\|A^{*}\|_{2}}\bigr{)}^{2}}\biggr{)}$
$\geq 1-\exp\biggr{(}-\frac{((1-p)\lambda T)^{2}}{16T\bigr{(}\frac{\bar{C}\sigma_{w}}{1-\|A^{*}\|_{2}}\bigr{)}^{2}}\biggr{)}$ (30)
for all $i,j\in\{1,\dots,d\}.$ The same lower bound holds for $\sum_{t=0}^{T-1}z_{t}^{ij-}$. In light of the union bound, both (14a) and (14b) hold for all $i,j\in\{1,\dots,d\}$ with probability at least $1-\delta$ if (30) is lower-bounded by $1-\frac{\delta}{2d^{2}}$. This provides the sufficient conditions for the exact recovery with probability at least $1-\delta$ due to Theorem 1. Thus, it suffices to require
$\hskip 28.45274pt1-\exp\biggr{(}-\frac{((1-p)\lambda T)^{2}}{16T\bigr{(}\frac{\bar{C}\sigma_{w}}{1-\|A^{*}\|_{2}}\bigr{)}^{2}}\biggr{)}\geq 1-\frac{\delta}{2d^{2}}$
$\Longleftrightarrow~{}\frac{((1-p)\lambda T)^{2}}{16T\bigr{(}\frac{\bar{C}\sigma_{w}}{1-\|A^{*}\|_{2}}\bigr{)}^{2}}\geq\log\Bigr{(}\frac{2d^{2}}{\delta}\Bigr{)}$
$\Longleftrightarrow~{}T\geq\biggr{(}\frac{4\bar{C}\sigma_{w}}{(1-p)\lambda(1-\|A^{*}\|_{2})}\biggr{)}^{2}\cdot\log\Bigr{(}\frac{2d^{2}}{\delta}\Bigr{)}.$
This completes the proof. ∎
Remark 2 (Uniqueness)
In general, multiple solutions can be attained from the minimization problem (11). In such a case, $A^{*}$ may not be the unique solution even after a sufficient time. The uniqueness is violated if the optimality condition is satisfied for some matrix other than $A^{*}$, which can be written as $A^{*}+\tilde{A}$ for some $\mathbb{R}^{d\times d}$. From (15), the solution $A^{*}$ is not unique if
$0\in\sum_{t=0}^{T-1}\partial\|w_{t}-\tilde{A}x_{t}\|_{1}\cdot x_{t}^{T}$ (31)
for some $\tilde{A}\neq 0$. Note that (16) and (31) are the same conditions if all $x_{t}$’s are zero, but this is already prevented by adopting Assumption 4. Furthermore, if the state trajectory $x_{0},\dots,x_{T-1}$ sufficiently spans the state space, which is determined by the rank of true matrix $A^{*}$ and the richness of the disturbance data, we can justify the uniqueness of the solution. Similar analyses have been presented in [zhang2024exact].
IV-C Exact Recovery under Arbitrary Disturbances
In the previous subsection, we have discussed the exact recovery under Assumption 5, which requires an equal probability for the sign of the disturbance to be positive or negative. In real-world applications, requiring exactly the same positive and negative probabilities may be challenging. By adopting the proofs used in the previous case, we will now present a more practical situation without considering Assumption 5: if the probability of having an attack at a given time is less than 0.5, it turns out that $l_{1}$-norm estimator overcomes any adversarial noncentral disturbance and the exact recovery is achieved. We formally state the theorem below.
Theorem 3
Suppose that Assumptions 1, 2, and 4 hold. Also, assume that Assumption 3 holds with $p<0.5$. Given $\delta\in(0,1]$, the $l_{1}$-norm estimator defined in (11) achieves the exact recovery after
$T\geq O\biggr{(}\biggr{(}\frac{\sigma_{w}}{(1-2p)\lambda(1-\|A^{*}\|_{2})}\biggr{)}^{2}\biggr{)}\cdot\log\Bigr{(}\frac{2d^{2}}{\delta}\Bigr{)}$ (32)
with probability at least $1-\delta$.
Proof:
In the scenario of $\mathbb{P}(w_{t}^{i}=0~{}|~{}\mathcal{F}_{t})=p<0.5$, for all $i,j\in\{1,\dots,d\}$, we have
$\mathbb{E}\Bigr{[}$ $\sum_{t=0}^{T-1}z_{t}^{ij+}\Bigr{]}$
$=\sum_{t=0}^{T-1}\mathbb{E}\Bigr{[}|x_{t}^{j}|\cdot\mathbb{P}(w_{t}^{i}=0~{}|~{}\mathcal{F}_{t})$
$\hskip 56.9055pt+x_{t}^{j}\cdot[\mathbb{P}(w_{t}^{i}<0~{}|~{}\mathcal{F}_{t})-\mathbb{P}(w_{t}^{i}>0~{}|~{}\mathcal{F}_{t})]\Bigr{]}$
$\geq\sum_{t=0}^{T-1}\biggr{[}(1-p)\mathbb{E}\bigr{[}|x_{t}^{j}|\bigr{]}-p\cdot\mathbb{E}\bigr{[}|x_{t}^{j}|\bigr{]}\biggr{]}\geq(1-2p)\lambda T>0,$
where the first equality is due to (22) in Lemma 3 and the first inequality comes from
$\bigr{|}\mathbb{P}(w_{t}^{i}<0~{}|~{}\mathcal{F}_{t})-\mathbb{P}(w_{t}^{i}>0~{}|~{}\mathcal{F}_{t})\bigr{|}\leq p,$
and the last inequality comes from $p<0.5$. Similarly, $\mathbb{E}[\sum_{t=0}^{T-1}z_{t}^{ij-}]$ has the same lower bound on the expectation. Since the expectations are positive, one can apply the proof of Lemma 4 to arrive at (32). Note that $1-p$ in (29) is simply replaced by $1-2p$. ∎
Remark 3 (Connection between Theorems 2 and 3)
Recall that Assumption 3 assumes $\mathbb{P}(w_{t}^{i}=0~{}|~{}\mathcal{F}_{t})>0$. In Theorem 2, Assumption 5 requires
$\mathbb{P}(w_{t}^{i}<0~{}|~{}\mathcal{F}_{t})=\mathbb{P}(w_{t}^{i}>0~{}|~{}\mathcal{F}_{t})=\frac{1-\mathbb{P}(w_{t}^{i}=0~{}|~{}\mathcal{F}_{t})}{2},$
which is strictly less than 0.5 for all $i=1,\dots,d$ due to Assumption 3. Now, suppose that every negative $w_{t}^{i}$ is infinitesimal. Then, the probability that $w_{t}^{i}$ is adequately far from zero is equivalent to the probability of positive attack which is less than 0.5, and therefore this approximately satisfies the assumptions of Theorem 3. This implies that if nonzero attacks occur with probability less than 0.5, the $l_{1}$-norm estimator can correctly recover the true system even when those attacks are always positive, which represents the worst-case scenario where the system may easily be deceived towards a positive bias. This surprising result shown by Theorem 3 serves as a special case of Theorem 2.
V Numerical Experiments
To be able to effectively demonstrate the results of this paper, we will provide two examples in this section.
Example 1
This example shows that under Assumption 5, the $l_{1}$-norm estimator successfully recovers the true matrix, which supports Theorem 2. We generate a random matrix whose operator norm is 0.6. Whenever the attack times happen with some probability $p<1$, each coordinate $w_{t}^{i}$ of the disturbance $w_{t}$ at time $t$ is selected by the adversary to be $-\text{sgn}(x_{t}^{i})\cdot\gamma$, where $\gamma$ follows $\text{Uniform}[-3,-1]$ with probability 0.5 and $\text{Uniform}[10,20]$ with probability 0.5. The disturbances satisfy the symmetry assumption given in (20), but indeed have a nonzero mean.
(a) $p=0.7,d=10$
(b) $p=0.7,d=20$
(c) $p=0.75,d=10$
(d) $p=0.75,d=20$
(e) $p=0.8,d=10$
(f) $p=0.8,d=20$
Figure 1: $l_{1}$-norm estimator vs. other estimators in Example 1.
In Figure 1, we show the efficacy of the $l_{1}$-norm estimator compared to the least-squares and the $l_{2}$-norm estimator brought up in Section I. We report the error based on the Frobenius norm of the difference between the estimates and the true matrix. Note that all three estimators involve convex optimization, and therefore they can be efficiently solved by standard optimization solvers. We present various scenarios based on the attack probability $p$ and the dimension $d$. While the OLS and $l_{2}$-norm estimator behave in a similar pattern, with a plateau in the error and failure to identify the correct matrix, the $l_{1}$-norm estimator achieves the exact recovery for all scenarios. By comparing the figures with the same dimension, one can easily observe that as $p$ increases, indicating more frequent attacks, the time required for the exact recovery also increases. Similarly, by comparing the figures with the same $p$, it can be seen that a larger dimension leads to a slower exact recovery. These findings illustrate the recovery time of $O\bigr{(}\frac{\log 2d^{2}}{(1-p)^{2}}\bigr{)}$ stated in Theorem 2.
(a) $p=0.45,d=10$
(b) $p=0.47,d=10$
(c) $p=0.48,d=10$
(d) $p=0.5,d=10$
Figure 2: $l_{1}$-norm estimator vs. other estimators in Example 2.
Example 2
For this example, we generated a random matrix whose operator norm is 0.95. We compare the $l_{1}$-norm estimator with the other estimators in the case of completely arbitrary disturbances with the attack probability $p<0.5$, which supports Theorem 3. We choose each coordinate $w_{t}^{i}$ of the disturbance $w_{t}$ to be Gaussian with a mean of $100\cdot(\text{sgn}(x_{t}^{i})+2)$ and a variance of 5 when the system is under attack. This adversarial disturbance implies that whenever the attack times happen, the disturbance is almost always positive, so as to mislead the estimators towards positive bias. However, Figure 2 shows that the $l_{1}$-estimator successfully identify the matrix when $p$ is at most 0.47, while the other estimators again fails to improve beyond a certain level of error. Figures 2(a) and 2(b) imply that a slight change in $p$ leads to a considerable difference in the exact recovery time. This is because, unlike in Example 1, the recovery time is modified to $O\bigr{(}\frac{\log 2d^{2}}{(1-2p)^{2}}\bigr{)}$, which is very sensitive near $p=0.5$. Figure 2(c) demonstrates this sensitivity as $p$ approaches 0.5, showing that the error alternates between zero and positive values. Finally, when $p$ is exactly 0.5, Figure 2(d) indicates that the $l_{1}$-norm estimator has a plateau in error, similar to that of the other estimators. This aligns with Theorem 3, which only holds for $p<0.5$. This observation identifies the critical number 0.5 when the $l_{1}$-norm estimator ultimately recovers the true matrix even in the presence of arbitrarily large noncentral attacks.
VI Conclusion
In this paper, we study the capability of the $l_{1}$-norm estimator to exactly identify the true matrix in the linear system identification problem, where the system suffers from adversarial noncentral disturbances. We show that the true matrix is exactly recovered under a symmetry assumption and the attack probability $p$ being less than 1. Furthermore, if $p<0.5$, the $l_{1}$-norm estimator prevails against any adversarial noncentral disturbances and achieves the exact recovery. This is the first result in the literature showing the possibility of accurate learning of systems under correlated, adversarial, and nonzero-mean disturbances.
[…]